# GCP: Guarded Collaborative Perception with Spatial-Temporal Aware Malicious Agent Detection

### v1

@@ V-C Qualitative Results
Visualization of BAC attack process. As shown in Fig. 6, the complete pipeline of BAC attack on V2X-Sim dataset is visualized. First, the malicious agent analyzes the detection results using only the victim vehicle’s local perception (Fig. 6(a)) to identify its inherent blind spots and perception limitations. Then, by comparing with the CP results (Fig. 6(b)), the attacker can identify regions where the victim heavily relies on collaborative messages for object detection. This differential analysis enables the malicious agent to generate an initial BAC seed map (Fig. 6(c)) that highlights these perception-dependent areas. Finally, through the proposed blind region segmentation algorithm, the attacker obtains a refined BAC confidence mask (Fig. 6(d)) that precisely delineates the victim’s vulnerable regions, providing targeted guidance for adversarial perturbation generation.
@@
Visualization of collaborative 3D detection. As shown in Fig. 7, we visualize the detection performance under different attack and defense scenarios on V2X-Sim dataset. While conventional attacks (PGD, C&W, BIM) generate obvious perturbations that baseline methods can easily detect, our BAC attack demonstrates superior stealthiness by optimizing output-space perturbations, specifically in blind regions. This makes it particularly challenging for existing methods like ROBOSAC and MADE to defend against, as evidenced by their significant performance degradation. In contrast, our proposed GCP framework maintains robust detection performance through effective spatial-temporal consistency verification, achieving results that approach the upper bound of CP. The visualization clearly shows that GCP successfully preserves accurate object detection even under sophisticated BAC attacks, while baseline methods struggle to maintain reliable performance.
@@
Visualization of BEV flow reconstruction loss distribution. Fig. 8 illustrates the distribution of BEV flow reconstruction losses across different agents. Normal agents (Agent 3 and 4) exhibit consistently low losses (mean: 0.0101 and 0.0091) with minimal variance and only one unmatched flow each, indicating stable temporal consistency. In contrast, agents under BAC attack (Agent 0 and 2) show significantly higher maximum losses (0.4219 and 0.8636) with greater variance and more unmatched flows (2 each), revealing temporal inconsistencies introduced by the attack. While some flows from malicious agents maintain low losses due to the stealthy nature of BAC attacks, the combination of high-loss outliers and increased unmatched flows provides reliable indicators for attack detection.
@@ VI Conclusion
In this paper, we have devised a novel blind area confusion (BAC) attack to show that collaborative perception (CP) systems are vulnerable to malicious attacks even with existing outlier-based CP defense mechanisms. The key innovation of the BAC attack lies in the blind region segmentation-based local perturbation optimization. To counter such attacks, we have proposed our GCP, a robust CP defense framework utilizing spatial and temporal contextual information through confidence-scaled spatial concordance loss and LSTM-AE-based temporal BEV flow reconstruction. These components are integrated via a spatial-temporal Benjamini-Hochberg test to generate reliable anomaly scores for malicious agent detection. Extensive experiments have demonstrated the superior robustness of our framework against various attacks while maintaining high detection performance. The effectiveness of our approach across different attack scenarios and perturbation settings highlights the potential of spatial-temporal analysis to guard CP systems. Looking forward, our work provides valuable insights into developing secure and reliable CP systems in autonomous driving, paving the way to robust multi-agent collaboration in real-world applications.
@@ Appendix A KF-based BEV Flow Interpolation
where $\mathbf{I}_{4}$ represents $4\times 4$ identity matrix, $\mathbf{0}_{4}$ is $4\times 4$ zero matrix, and $\Delta t$ is the time difference between frame $k$ and frame $k-1$. Based on the state transition matrix above, there are two stages for KF, namely, the prediction stage and the update stage. In the prediction stage, both the state and the error covariance are predicted, given as:

### v2

@@ V-D Qualitative Results
Visualization of Attack Pipeline. Fig. 7 illustrates the complete pipeline of the BAC attack on the V2X-Sim dataset. The malicious agent analyzes detection results from the victim’s local perception (Fig. 7(a)) to identify blind spots. By comparing with CP results (Fig. 7(b)), it pinpoints regions where the victim relies heavily on collaborative messages. This differential analysis yields an initial seed map (Fig. 7(c)), which is refined into a BAC confidence mask (Fig. 7(d)) via blind region segmentation to guide targeted perturbations.
@@
Visualization of Defense Performance. We visualize the collaborative 3D detection performance in Fig. 5. While conventional attacks (PGD, C&W, BIM) generate obvious perturbations that are easily detected, our BAC attack optimizes output-space perturbations in blind regions, demonstrating superior stealthiness and causing significant degradation in baselines like ROBOSAC and MADE. In contrast, GCP maintains robust detection through spatial-temporal consistency verification, achieving results comparable to the upper bound.
@@
Visualization of BEV Flow Reconstruction Loss. Fig. 8 analyzes the BEV flow reconstruction loss distribution. Normal agents (Agent 3 and 4) show consistently low losses (mean $\approx$ 0.01) and stable temporal consistency. In contrast, agents under BAC attack (Agent 0 and 2) exhibit significantly higher maximum losses (>> 0.4) and more unmatched flows. These high-loss outliers and temporal inconsistencies serve as reliable indicators for detecting stealthy attacks.
Visualization of Statistical Validity. We validate the statistical assumptions on a small calibration set comprising 389 agent-frame samples from V2X-Sim Scene 8. The BH procedure requires the test statistics to satisfy Positive Regression Dependence on a Subset (PRDS). As shown in Fig. 9, we observe consistently non-negative cross-agent correlations ($r\in[0.02,0.51]$) due to shared field-of-views, alongside positive temporal autocorrelation from physical continuity. This positive dependence structure satisfies the PRDS condition, ensuring the validity of FDR control. Fig. 9(d) further confirms that our chosen operating point ($\alpha_{bh}=0.36$) achieves 100% attack recall while maintaining conservative FDR control, effectively balancing safety and benign performance.
V-E Limitation and Future Work
Currently, we focus on defending against untargeted adversarial attacks without temporal optimization. While sophisticated “slow-drift” attacks could potentially evade detection by generating temporally consistent long-term trajectories, such methods are currently unexplored in the context of real-time multi-agent systems due to the optimization difficulities. Therefore, investigating the feasibility of such stealthy threats and developing corresponding defense mechanisms remains an important direction for future research.
@@ VI Conclusion
In this paper, we have devised a novel blind area confusion (BAC) attack to show that collaborative perception (CP) systems are vulnerable to malicious attacks even with existing outlier-based CP defense mechanisms. The key innovation of the BAC attack lies in the blind region segmentation-based local perturbation optimization. To counter such attacks, we have proposed our GCP, a robust CP defense framework utilizing spatial and temporal contextual information through confidence-scaled spatial concordance loss and LSTM-AE-based temporal BEV flow reconstruction. These components are integrated via a spatial-temporal Benjamini-Hochberg test to generate reliable anomaly scores for malicious agent detection. Extensive experiments demonstrate GCP’s superior robustness against diverse attacks while preserving benign performance. These results validate the efficacy of spatial-temporal analysis in guarding CP systems. We believe our work establishes a solid foundation for secure multi-agent collaboration in future autonomous driving applications.
@@ Appendix A KF-based BEV Flow Interpolation
where $\mathbf{I}_{4}$ represents $4\times 4$ identity matrix, $\mathbf{0}_{4}$ is $4\times 4$ zero matrix, and $\Delta t$ is the time difference between frame $k$ and frame k1k-1. Based on the state transition matrix above, there are two stages for KF, namely, the prediction stage and the update stage. In the prediction stage, both the state and the error covariance are predicted, given as: